explain step by step not getting the correct numbers

10.00 points John Jones owns and manages a cafe in Collegetown whose annual revenue is $5,000. Annual expenses are as fallows: Expenae abor ood and dink ectnc ty ehicle lase ent nterest on Inan tor equipment だ1(1 a) Suppose John could earn $1,100 per year as a recycler of aluminum cans. However, he prefers to run the cafe and is vwilling to pay $275 per year to run the cafèe rather than to recycle. Is the cafe making an economic profitr , the café is making an economic kos of -350 per year Should John stay in the cafe business? No, he should not stay in the cefé business. b) Suppose John had not had to get a $10,000 loan at an annual interest rate of 10 percent to buy equipment, but instead had invested $10,000 of his own money in equipment. Calculate Johns annual accounting profit. $10rbl c) Suppose John could eam $1,000 a year as a recycler, and he likes recycling just as well as running the cafe How much additional revenue would the café have to collect each year to eam a normal pront? 2501
Answer #1

Economic loss = revenue - accounting cost ( labour + food and drink + electricity + vehicle lease + rent + interest on loan ) - economic cost ( earning from recycling ) - payment to run cafe

=5000 - (2000 + 500 + 100 + 150 + 500 + 1000 ) -1100 - 275

= -625

economic cost includes the opportunity cost of next best alternative (in this case earning from recycling operations + payment to run cafe ) + accounting cost ( payment made in cash )

No, the cafe is not making economic profit .

loss = $625

b)Accounting profit = revenue from operations - accounting cost (payment made ).

= 5000 - accounting cost ( labour + food and drink + electricity + vehicle lease + rent ). ( since the equipment is bought from own money so no interest will be paid on equipment so therefore interest will not be included in the cost )

= 5000 - (2000 + 500 + 100 + 150 + 500)

= 1750

c) Normal profit means 0 profit as normal profit covers both accounting cost and economic cost .

also as he prefers both activities equally ,

so revenue = 5000

accounting cost = 4250

economic cost (earning from recycling ) = $1000

loss = 5000 - 5250

= 250

so he need to earn $250 more to earn normal profit .